# Ettin: Analyzing Encoders vs Decoders Using the Same Architecture and Data

### v1

@@ 1 Introduction
Part of this gap is due to the sentiment within the community that decoders can be adapted for use in tasks that were once predominantly encoder-focused (e.g. classification, embeddings), especially as they can often be used in a zero-shot fashion (i.e. without fine-tuning). As decoder models are more studied, more over-trained (Hoffmann et al., 2022), and are generally larger, they are claiming the top of the leaderboards for previously encoder-centric tasks (Enevoldsen et al., 2025).
Many works have challenged this assumption by comparing encoder-only and decoder-only models of roughly the same sizes (Ethayarajh, 2019; Charpentier & Samuel, 2024; Harrag et al., 2021). However, these analyses have to be done with incomparable models: using different architectures, different pre-training data, different learning schedules, etc.
@@
We find that, like previous work, encoders excel at classification and retrieval while decoders excel at generative tasks. However, we go beyond previous work to examine the increasingly common setting (BehnamGhader et al., 2024) where decoder-models are continued trained as encoders (i.e. cross-objective training). We show results for this cross-objective training in both directions: training encoders for casual language modeling (CLM) and decoders with masked language modeling (MLM). We find that despite continued training for much longer than previous work (50B tokens) these models do not surpass those that started with this objective, i.e. a 400M encoder outperforms a 1B decoder continue-trained with MLM on MNLI, and vice versa for generative tasks.
@@ 2 Related Work
Previous work has compared encoder-only and decoder-only models on a wide assortment of tasks. For example Charpentier & Samuel (2024) compare DeBERTa and GPT-2 in similar sizes. Other work (Yang et al., 2023; Qu et al., 2020; Zheng et al., 2021; Rehana et al., 2023) compares them on downstream tasks.
However, all of these comparisons have the same underlying limitation: the models they are comparing have different numbers of parameters, different architectures, different training recipes, and different pre-training data. Although some work has attempted to address this (Charpentier & Samuel, 2024; Gisserot-Boukhlef et al., 2025), they have only done so in limited settings with very small amounts of pre-training data. In contrast, we train SOTA models, allowing for an exact comparison.
@@
Although we cannot cover it all here, there have been attempts to use bidirectional attention for standard decoder usage. This includes prefix LM attention (Artetxe et al., 2022; Chowdhery et al., 2023; Du et al., 2021) and other mixed training such as BiTune (Kopiczko et al., 2024). However, most modern LM training still favors pure CLM. [^4] The the best of our knowledge, as much of the details of the best LMs now goes unpublished.
@@ 3.3 Training Recipe
We use the same general process described by open-data models (which was followed by ModernBERT) for training both encoder-only and decoder-only models – with a few specific changes for the encoder architecture (i.e. masking ratio). In summary, we include three general phases: base pre-training, mid-training/context extension, and decay. See Table 2 for the precise sources of training data in each phase. We use a trapezoidal learning rate scheduler, with general hyperparameters shown in Appendix C and size-dependent hyperparameters in Table 1.
@@ 3.6 Cross-Objective Training
As encoder models have gone out of popularity, decoder models have increased in size (both parameters and pre-training data). Thus, these newer decoder models are typically trained for much longer than previous encoder models (i.e. BERT). Due to this it has become common to adapt these larger decoder models to what was previously encoder-centric tasks (Zhang et al., 2025). With paired encoder and decoder models, we can now answer the question of how effective this continued pre-training approach is and whether it is still worth training both types of models. We call this cross-objective training: taking the final model and continue pre-training it on the reverse objective. [^8] Following BehnamGhader et al. (2024), we do not use MLM but rather use MNTP, that is, the masked token is predicted using the hidden state of the previous token to better align with CLM.
We train for 50B tokens with the reverse objective, which is far more than previous work has attempted (BehnamGhader et al., 2024), which is usually around 10B tokens. Although the ratio of pre-training and cross-objective training is unbalanced, this mimics the realistic setting where the adaptation is often done very small amounts of data comparatively. We do this cross-objective training on the highest quality data we have available, which was used in the last decay phase. [^9] We note that this means it repeating this data twice, however, as shown by previous work (Muennighoff et al., 2023) two repetitions on high quality data has no adverse effects. We use a new trapezoidal learning rate schedule with 3B tokens of warmup and 10B tokens of decay. [^10] For the 1B model this is scaled by 1/3 again due to compute availability. Thus, by the end we have an encoder-from-decoder (i.e. a decoder further pretrained with MLM [^11] We use a 15% masking rate for the encoder-from-decoder as to maintain a middle ground masking ratio.) and a decoder-from-encoder (i.e. an encoder further pre-trained with CLM).
@@ 4.1 Individual Evaluations
We use two baselines for each size type: extra extra small (XXS) BERT-mini and TinyBERT, extra small (XS) models MiniLM L12 and BERT-small, small (S) models DistilBERT and DistilRoBERTa, base (B) models BERT and ModernBERT, large (L) models BERT-large and ModernBERT-large, and an extra large (XL) model DeBERTA v2 XL. [^12] We also ran experiments with DeBERTa XXL as shown in the appendix. However, due to the size and slowness of the architecture we could not do an comparable grid search. Our results in the appendix are after 300 days of H100s hours, but still did not complete the full sweep. Furthermore, as DeBERTa XXL is ¿ 1.5B we exclude it as it is significantly larger than 1B (i.e. ¿ 50% larger).
We evaluate on various encoder tasks, including GLUE (Wang et al., 2018), MTEB v2 English (Enevoldsen et al., 2025), MDLR for long context (Chen et al., 2024), and CodeSearchNet for code evaluation (Husain et al., 2019). We use the same evaluation setup as ModernBERT for the evaluation for an equal comparison (see Appendix E for hyperparameter details).
@@
Table 4: Performance comparison of decoder-only models across tasks, organized by size categories. We see that Ettin decoders compare favorably, matching or exceeding the previous open-data SOTA. Task names in order are ARC, Hellaswag, LAMBADA, OpenBookQA, Social IQA, TrivialQA, Winogrande, and Winograd Schema Challenge.
@@
We evaluate on a wide range of tasks using the Eleuther AI harness (Gao et al., 2024a) (see Appendix B for details), consolidating tasks used in the Pythia and SmolLM papers including: the ARC Challenge (ARC) Clark et al. (2018), HellaSwag (HS) (Zellers et al., 2019), LAMBADA (LMB) (Paperno et al., 2016), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social IQA (SIQA) (Sap et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), Winogrande (WG) (Sakaguchi et al., 2021), and the Winograd Schema Challenge (WSC) (Levesque et al., 2012).
@@
Figure 1: Encoder vs decoder comparison across model size. In most cases, models in the preferred architecture (e.g. encoders in MNLI) do better than the opposite architecture even with an order of magnitude greater size, e.g. a 400M decoder outperforming the 1B encoder. Notably, in generative tasks, decoders-from-encoders scale poorly with size, matching performance at small sizes but seeing a wide gap as model size grows. A table version of this plot can be found in Table 8.
@@ 5 Case study: Gender Bias
Due to the Ettin suite's open-pretraining data, we can also analyze other aspects of learning across pre-training objectives. As one example of future research that can be done, we analyze which objective learns more biased gender representations.
@@
As both models saw the exact same training data, we see that the MLM objective encourages the model to choose more neutral pronouns over female pronouns. However, male gender bias seems strong in both models, if slightly higher for decoders. This shows just one type of training-data based analysis that can be done and we leave others to future work.
@@ 6 Discussion
This has several implications for those using models for classification or retrieval: currently the top models on leaderboards like MTEB are 7B+. However, based on our experiments, it is likely a 3B encoder model would outperform it. The lack of large encoder-only size models means that approaches that continue pre-train decoders will likely outperform all other options (as is currently seen on the leaderboards). In the small scale regime (1B or less) it easier to train more ``niche" encoder models for classification and retrieval and based on our results it seems likely encoders will continue to outperform all others in their size range.
Our results also suggest that encoders and decoders learn differently in other aspects as well, such as gender bias. Although this is just one example, we look forward to future research that discovers other differences and provide our artifacts to enable such work.
@@ Appendix D Model Sizes
Appendix E Encoder Evaluation Sweep Parameters
@@
[…]
@@
[…]
As Ettin models have been trained with instructions during pre-training, it is likely they are also more capable for instruction-based retrieval (Shao et al., 2025; Weller et al., 2024, 2025), however, we leave that for future work.

### v2

# Seq vs Seq: An Open Suite of Paired Encoders and Decoders

@@ 1 Introduction
Part of this gap is due to the sentiment within the community that decoders can be adapted for use in tasks that were once predominantly encoder-focused (e.g. classification, embeddings), especially as they can often be used in a zero-shot fashion (i.e. without fine-tuning) (BehnamGhader et al., 2024). As decoder models are more studied, more over-trained (Hoffmann et al., 2022), and are generally larger, they are now claiming the top spots of leaderboards for previously encoder-centric tasks (Enevoldsen et al., 2025).
Many works have challenged this assumption by comparing encoder-only and decoder-only models of roughly the same sizes (Ethayarajh, 2019; Charpentier and Samuel, 2024; Harrag et al., 2021). However, these analyses have to be done with incomparable models: using different architectures, different pre-training data, different learning schedules, etc.
@@
We find that, like previous work, encoders excel at classification and retrieval while decoders excel at generative tasks. However, we go beyond previous work to examine the increasingly common setting (BehnamGhader et al., 2024) where decoder-models are continued trained as encoders (i.e. cross-objective training). We show results for this cross-objective training in both directions: training encoders for causal language modeling (CLM) and decoders with masked language modeling (MLM). We find that despite continued training for much longer than previous work (50B tokens) these models do not surpass those that started with this objective, i.e. a 400M encoder outperforms a 1B decoder continue-trained with MLM on MNLI, and vice versa for generative tasks.
@@ 2 Related Work
Previous work has compared encoder-only and decoder-only models on a wide assortment of tasks. For example Charpentier and Samuel (2024) compare DeBERTa and GPT-2 in similar sizes. Other work (Yang et al., 2023; Qu et al., 2020; Zheng et al., 2021; Rehana et al., 2023; Nielsen et al., 2024) compares them on downstream tasks.
However, all of these comparisons have the same underlying limitation: the models they are comparing have different numbers of parameters, different architectures, different training recipes, and different pre-training data. Although some work has attempted to address this (Charpentier and Samuel, 2024; Gisserot-Boukhlef et al., 2025), they have only done so in limited settings with very small amounts of pre-training data. In contrast, we train SOTA models, allowing for an exact comparison.
@@
Although we cannot cover it all here, there have been attempts to use bidirectional attention for standard decoder usage. This includes prefix LM attention (Artetxe et al., 2022; Chowdhery et al., 2023; Du et al., 2021) and other mixed training such as BiTune (Kopiczko et al., 2024). However, most modern LM training still favors pure CLM. [^4] To the best of our knowledge, as much of the details of the best LMs now goes unpublished.
@@ 3.3 Training Recipe
We use the same general process described by open-data models (which was followed by ModernBERT) for training both encoder-only and decoder-only models – with a few specific changes for the encoder architecture (i.e. masking ratio). In summary, we include three general phases: base pre-training, mid-training/context extension, and decay. See Table 2 for the precise sources of training data in each phase. We use a trapezoidal learning rate scheduler, with general hyperparameters shown in Appendix C and size-dependent hyperparameters in Table 1. For compute details see Appendix E.
@@ 3.5 Cross-Objective Training
As encoder models have gone out of popularity, decoder models have increased in size (both parameters and pre-training data). Thus, these newer decoder models are typically trained for much longer than previous encoder models (i.e. BERT). Due to this it has become common to adapt these larger decoder models to what were previously encoder-centric tasks (Zhang et al., 2025). With paired encoder and decoder models, we can now answer the question of how effective this continued pre-training approach is and whether it is still worth training both types of models. We call this cross-objective training: taking the final model and continue pre-training it on the reverse objective. Following BehnamGhader et al. (2024), we do not use MLM but rather use MNTP, that is, the masked token is predicted using the hidden state of the previous token to better align with CLM.
We train for 50B tokens with the reverse objective, which is far more than previous work has attempted e.g. LLM2Vec (BehnamGhader et al., 2024), which is around 10B tokens. Although the ratio of pre-training and cross-objective training is unbalanced, this mimics the realistic setting where the adaptation is done with very small amounts of data comparatively. We do this cross-objective training on the highest quality data we have available, which was used in the last decay phase. [^8] We note that this means it repeating this data twice, however, as shown by previous work (Muennighoff et al., 2023) two repetitions on high quality data has no adverse effects. We use a new trapezoidal learning rate schedule with 3B tokens of warmup and 10B tokens of decay. [^9] For the 1B model this is scaled by 1/3 again due to compute availability. Thus, by the end we have an encoder-from-decoder (i.e. a decoder further pretrained with MNTP similar to LLM2Vec [^10] We use a 15% masking rate for the encoder-from-decoder as to maintain a middle ground masking ratio.) and a decoder-from-encoder (i.e. an encoder further pre-trained with CLM).
@@ 4.1 Individual Evaluations
We use two baselines for each size type: extra extra small (XXS) BERT-mini and TinyBERT, extra small (XS) models MiniLM L12 and BERT-small, small (S) models DistilBERT and DistilRoBERTa, base (B) models BERT and ModernBERT, large (L) models BERT-large and ModernBERT-large, and an extra large (XL) model DeBERTA v2 XL. [^11] We also ran experiments with DeBERTa XXL as shown in the appendix. However, due to the size and slowness of the architecture we could not do a comparable grid search. Our results in the appendix are after 300 days of H100s hours, but still did not complete the full sweep. Furthermore, as DeBERTa XXL is ¿ 1.5B we exclude it as it is significantly larger than 1B (i.e. ¿ 50% larger).
We evaluate on various encoder tasks, including GLUE (Wang et al., 2018), MTEB v2 English (Enevoldsen et al., 2025), MDLR for long context (Chen et al., 2024), and CodeSearchNet for code evaluation (Husain et al., 2019). We use the same evaluation setup as ModernBERT for the evaluation for an equal comparison (see Appendix F for hyperparameter details).
@@
Table 4: Performance comparison of decoder-only models across tasks, organized by size categories. We see that Ettin decoders compare favorably, matching or exceeding the previous open-data SOTA. Task names in order are ARC, Hellaswag, LAMBADA, OpenBookQA, Social IQA, TriviaQA, Winogrande, and Winograd Schema Challenge.
@@
We evaluate on a wide range of tasks using the Eleuther AI harness (Gao et al., 2024a) (see Appendix B for details), consolidating tasks used in the Pythia and SmolLM papers including: the ARC Challenge (ARC) Clark et al. (2018), HellaSwag (HS) (Zellers et al., 2019), LAMBADA (LMB) (Paperno et al., 2016), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social IQA (SIQA) (Sap et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), Winogrande (WG) (Sakaguchi et al., 2021), and the Winograd Schema Challenge (WSC) (Levesque et al., 2012). Note all tasks use zero-shot closed-book evaluations.
@@
Figure 1: Encoder vs decoder comparison across model size. Generally models in the preferred architecture (e.g. encoders in MNLI) do better than the opposite architecture even with an order of magnitude greater size, e.g. a 400M decoder outperforming the 1B encoder. Notably, in generative tasks, decoders-from-encoders scale poorly with size.
@@ 4.2 Encoders vs Decoders
Although most of our models are too small to have signal on more difficult, but standard LM tasks like MMLU and GSM8k, we evaluate these for just the 1B sized models in Table 5. We find similar results to the previous table, illustrating that the Decoder-From-Encoder significantly outperforms on MMLU classification but is significantly worse on the generative GSM8k.
Model MMLU CS GSM8k
Decoder 27.0 32.0
Dec-From-Enc 37.0 18.9
Table 5: Harder benchmark results for 1B models on reasoning and classification tasks.
@@ 5 Case study: Gender Bias
Due to the Ettin suite's open-pretraining data, we can also analyze other aspects of learning across pre-training objectives. As one example, we analyze gender representations for bias.
@@
As both models had the same training data, we find that the MLM objective leads the model to choose more neutral pronouns over female pronouns. However, male gender bias seems strong in both models, if slightly higher for decoders. Thus, this is one example of the analysis enabled by our data; we leave others to future work.
@@ 6 Discussion
This has several implications for those using models for classification or retrieval: currently the top models on leaderboards like MTEB are 7B+. However, based on our experiments, it is likely a 3B encoder model would outperform it. But, the lack of large encoder-only models means that approaches that continue pre-train decoders using MLM will likely outperform all other options (as is currently seen on the leaderboards). In the small scale regime (1B or less) where it is easier to train more ``niche" encoder models for classification/retrieval, our results indicate that encoders will continue to outperform all others in their size range (and even ones above it).
Our results also suggest that encoders and decoders learn differently in other aspects as well, such as gender bias. Although this is just one example, we look forward to future research that discovers other differences. Overall, our artifacts allow for a range of new analyses and pre-training research.
We note that concurrent work (Gisserot-Boukhlef et al., 2025) did a similar analysis and found that starting from CLM and doing continued MLM pre-training was better in nearly all cases. However, pre-training was done with only 100B tokens; thus it is likely that this is an artifact of CLM's comparative data efficiency in the smaller data regime (i.e. loss on every token compared to loss on X% of tokens for MLM). Our work shows that at scale, for SOTA models, encoders do greatly outperform encoders-from-decoders (as also shown by SOTA concurrent work Marone et al. (2025)).
@@ Appendix D Model Sizes
Appendix E Compute Configuration
Appendix F Encoder Evaluation Sweep Parameters
@@
We re-use ModernBERT's evaluation setup but slightly increase the learning rate sweep in order to better fit the smaller parameter models (which typically use higher LRs). As the best LRs chosen by BERT and ModernBERT were lower than these, it does not affect their scores. We sweep for learning rates over $\{1\text{e-}5,3\text{e-}5,5\text{e-}5,8\text{e-}5,1\text{e-}4\}$, weight decay values over $\{1\text{e-}6,5\text{e-}6,8\text{e-}6,1\text{e-}5\}$, and batch sizes over $\{16,32\}$. We also sweep over epochs $\{1,2,3,4\}$ if task $\in\{\text{mnli},\text{sst2},\text{rte}\}$, otherwise $\{2,5,10,12\}$. This was a total of 160 sweeps per model, of which we select the best score per task per model to report. We start from the best MNLI checkpoint for fine-tuning on RTE, STS-B, and MRPC, following ModernBERT.
@@
We sweep four LRs ($\{1\text{e-}4,3\text{e-}4,5\text{e-}4,7\text{e-}4\}$) on MS MARCO dev and choose the best performing one to evaluate on the other retrieval datasets. We use a new retrieval training script due to being unable to exactly reproduce ModernBERT's precise scores. While doing so, we also improve the training process for all models due to the use of more negatives in training, achieving higher scores than that in the ModernBERT paper (which was not trying to optimize scores, but shows generally that our training script is effective). We trained with an effective batch size of 1024 with 4 accumulation steps. For DeBERTa-v2 it diverged for all learning rates we tried. Thus, to get it to converge, we changed the warmup to 20% from 5% and lowered the learning rate to 1e-5.
As Ettin models have been trained with instructions during pre-training, it is likely they are also more capable for instruction-based retrieval (Shao et al., 2025; Weller et al., 2024; 2025), however, we leave that for future work.
Appendix G LLM Usage
LLMs were not used for paper writing, but were used for coding assistance and title brainstorming. All code was human verified.